a vertical u tube of uniform cross section contains mercury in both sides.a glycerine ( density 1.3 g/c.c) column of 10 cm is introduced into one arm.Oil of density 0.8 g/c.c is poured into the other arm until the oil and glycerine are in the same horizontal level; Find the length of oil column. density of mercury= 13.6g/c.c

Grade:

4 Answers

mohit yadav
11 years ago
3.1
SOURAV MISHRA
11 years ago

Let the length of the oil column be x cm. 

we can write

0.8*x*g + 13.6*(10 - x)*g = 1.3*10*g     where g is the acceleration due to gravity.

this gives x = 123/12.8 = 9.61 cm. 

the length of the oil column is 9.61 cm.

Koushik
6 years ago
The glycerin and oil + some amount of mercury should be at   
some equipressure line.   
Let height of oil is h then
1.3 *g*10=0.8*g*x+13.6*g*(10-x)
Therefore, x =9.6 cm
Kushagra Madhukar
4 years ago
Dear student,
Please find the solution to your problem below.
 
Let the length of the oil column be x cm.
we can write
0.8*x*g + 13.6*(10 - x)*g = 1.3*10*g     [where g is the acceleration due to gravity]
This gives x = 123/12.8 = 9.61 cm
 
Thanks and regards,
Kushagra